# Surface topological quantum criticality: Conformal manifolds and Discrete Strong Coupling Fixed Points

### v1

@@ IV.4 Symmetries of the interaction matrix $V_{ij}$
IV.5 Renormalization group study of the effective interacting model for a general $\mathcal{N}$: $2+\epsilon$ expansion
@@
So far, we have developed an effective interacting model for the TI surface. We are now ready to study the surface quantum criticality using the renormalization-group-equation (RGE) technique. In Eqn.LABEL:Lfrfmn, we have an interacting theory of $\mathcal{N}$ flavors of 2-component Dirac fermions in $2+1D$. Dimensional analysis suggests that the interaction term in Eqn.LABEL:Lfrfmn is marginal in $1+1$D dimensions. So in $(2+1)D$, the interaction term is irrelevant in the infrared (IR) limit if the interaction is weak. However, in the case of $\mathcal{N}=1$, it had been demonstrated that the interacting theory has an interacting UV fixed point [60, 61, 15, 62, 16, 17, 18]. Here, we study surface quantum criticality phenomena in a more general case with $\mathcal{N}>1$ surface Dirac cones and search for strong coupling conformal-field-theory fixed points that shall define possible universality classes of surface topological quantum criticality for a given $\mathcal{N}$.
@@
Note that by increasing $l$, we are moving towards the IR limit of the theory. The numerical constant $c^{2}=\frac{\pi}{8}$. $V_{ij}$ is the renormalized dimensionless interaction matrix element. The derivation details of Eqn.26 are moved to the appendix.B for brevity.
@@
Note that the $SO(\mathcal{N})$ symmetry of the RG equation in Eqn.27 is $emergent$ and not fundamental. The theory does not have flavor rotation symmetry. The RG equation has this symmetry because we limited our calculations to one-loop contributions. A higher-loop contribution explicitly breaks the flavor rotation symmetry, as evident from the two- or three-loop diagrams shown in Fig.6 (b), (c), and (d).
Table 2: Table showing the list of fixed points, the shape of the manifold in the parameter space, and their stabilities for the $\mathcal{N}=2$ flavor surface. Here $\phi\in[0,2\pi]$.
@@
$\hat{V}_{c}=\frac{\epsilon}{2}\hat{I}+\frac{\epsilon}{2}\cos\phi\,\sigma_{z}+\frac{\epsilon}{2}\sin\phi\,\sigma_{z}$ One relevant, one irrelevant and one marginal operator. $\hat{I}$ ring
@@
V Symmetries of the RGE and its dynamical solutions
Before we begin discussing the fixed points and their properties for different $\mathcal{N}$, here we present a fascinating observation from our study on the fixed points and the formation of conformal manifolds. Whether a set of fixed points forms a smooth manifold is conveniently and intuitively related to the dynamical symmetry of the RG equation and the invariant subgroup of the fixed point coupling matrix. In this section, we shall put forward the general argument. In the upcoming sections, where we solve for the fixed points of the RG Eqn.27 for $\mathcal{N}=2$ and $\mathcal{N}=3$ flavor surfaces, we shall show explicitly that this general argument is followed in all cases.
@@ V Symmetries of the RGE and its dynamical solutions
If the dimension of the coset space is greater than zero, a smooth conformal manifold is formed in the parameter space. This symmetry argument proves to be very useful in understanding the shape and dimensions of the conformal manifolds if it exists. We shall make explicit use of this theory to solve the fixed points of the RG equation (Eqn.27) in the case of $\mathcal{N}=3$ TI surface, where the parameter space is six-dimensional and a direct brute-force calculation is challenging.
Interestingly, we shall find in section.VIII that the symmetry of the dynamical equation is $not$ a necessary condition for the existence of conformal manifolds. When considering the 3-loop effects in the RG equation, we find that the $SO(\mathcal{N})$ symmetry is broken (see Fig.6). However, the conformal manifold persists, although in a distorted shape.
VI Conformal manifolds and phase boundary in the $\mathcal{N}=2$ flavor TI surface
@@
The main objective here is to study the isolated fixed points and conformal manifolds of the interacting theory in the three-dimensional parameter space ($D_{p}=3$). Once the fixed point manifolds are found, we intend to demonstrate how the conformal manifolds or isolated fixed points can be represented as a coset space $G/H$, as outlined in section.V. Following this, we shall look for the phase boundary that separates the gapless surface state from the superconducting state. Finally, we show that the infrared physics at the phase boundary can be conveniently studied via an effective Yukawa field theory that involves a single complex boson strongly interacting with two flavors of gapless fermions.
The interaction matrix $\hat{V}$ defined in Eqn.27 is a $2\times 2$ symmetric matrix. Thus, the RG equation possesses emergent $SO(2)$ symmetry. Given that $\hat{V}$ is a real symmetric matrix, we could re-parametrize the $\hat{V}$-matrix in the following simple way:
@@
where $\sigma_{x}$ and $\sigma_{z}$ are Pauli matrices defined in the parameter space. This decomposition makes sense because a real symmetric matrix of size $2$ has $3$ independent parameters, which is the case here. Physically, $V_{x}$ signifies the scattering of Cooper pairs between the surface fermions of different flavors. $V_{z}$ represents the difference in the intra-flavor interactions of surface fermions. Plugging this back to the RG equation in Eqn.27, we get,
@@
Following the general discussions about the $SO(\mathcal{N})$ symmetry of Eq.27, 28, one can easily show that the RG flow lines exhibit an axial $SO(2)$ rotation symmetry. This is explicitly evident from the invariance of the RGE(given in Eqn.32) under $\hat{R}(\chi)$, $SO(2)$ rotations about the $V_{0}$ axis in the $V_{0}-V_{x}-V_{y}$ representation introduced in Eq.31,
@@
Another important consequence of the $SO(2)$ invariance is that the scaling dimensions of the operators are the same for all points on the ring manifold (see Appendix.C).
VI.1 Fixed points Vs Conformal manifolds
@@
Figure 7: RG flows of the parameters $V_{0}$, $V_{z}$ and $V_{x}$ defined in Eqn.31. The red points denote the isolated fixed points, while the red-colored ring denotes the fixed point manifold. The blue line indicates the direction of an irrelevant operator corresponding to the $V_{x,c}=0,V_{z,c}=\frac{\epsilon}{2},V_{0,c}=\frac{\epsilon}{2}$ point on the ring manifold while the green line represents a relevant operator.
@@
$\hat{V}_{c}=\frac{\epsilon}{2}\hat{I}+\frac{\epsilon}{2}\cos\phi\,\sigma_{z}+\frac{\epsilon}{2}\sin\phi\,\sigma_{z},\,\phi\in[0,2\pi]$
@@
One can see that for the fixed point manifold, there is one irrelevant, one relevant, and one marginal operator. The marginal operator at a point on the ring manifold is directed tangential to it. The relevant operator is directed along the line connecting the origin to the given point in the manifold.
The RG flow diagram is shown in fig.7. Appendix C lists all the relevant, irrelevant, and marginal operators and their respective scaling dimensions at any point on the ring manifold. This ring manifold can be described as a conformal manifold of co-dimension $D_{co}=2$ in the three-dimensional parameter space.
VI.2 Conformal manifolds as a coset space $G/H$
Having found out all the fixed points from the RG equation, let us now see how the argument made in section V on the connection between the dynamical symmetries and the geometry of the coset manifolds can be applied to each of the cases. We show here explicitly that the geometry of coset space for each fixed point matches what we found above using brute calculations.
@@ VI.3 Phase boundary
VI.4 Effective Yukawa field theory for the phase boundary
@@
Let us revisit the expression of the interaction matrix $\hat{V}_{c}$ at the conformal manifold once again. From the table.2, it has the form, $\hat{V}_{c}=\frac{\epsilon}{2}\hat{I}+\frac{\epsilon}{2}\cos\phi\,\sigma_{z}+\frac{\epsilon}{2}\sin\phi\,\sigma_{z}$. After simple rearrangements, $\hat{V}_{c}$ can be written as a tensor product of a vector in 2-dimensional space with itself. Let $\textbf{n}=\sqrt{\epsilon}\left(\cos\theta,\,\,\,\sin\theta\right)$ where $\theta=\phi/2$ be the unit vector. Then, $\hat{V}_{c}$ can be expressed in terms of n as,
@@
where $m$ is the boson mass and $\lambda$ is the four-boson interaction term. Here, we introduced a new term $g_{i}$ instead of $n_{i}$ used in Eqn.48 because after adding the boson dynamics into the theory, the effective Yukawa theory in Eqn.49 is quantitatively different from the one in Eqn.48. As mentioned at the beginning of this subsection, the boson mass term can be considered effectively as that relevant interaction that must be tuned to zero to study the quantum critical properties of the theory.
@@
VII Conformal manifolds and phase boundary in the $\mathcal{N}=3$ flavor TI surface
@@
We constructed the effective 2D action governing the interaction in Eqns.21, LABEL:Lfrfmn. In section IVD, we derived the RGE by studying the interaction up to the one-loop level. The interaction matrix $\hat{V}$ in Eqn.27 is now a $3\times 3$ symmetric matrix, and the interaction parameter space is six-dimensional ($D_{p}=6$). Therefore, a brute-force solution would be challenging. Here is where the emergent $SO(3)$ symmetry of the RGE comes to the rescue. It suggests that if they exist, the fixed point matrices $\hat{V}_{c}$ could be distinguished by their eigenvalues(or traces). Therefore, we shall first look at the RG flow of the eigenvalues of the $\hat{V}$ matrix in section.VIIA.
Once we find the fixed points in terms of the eigenvalues, we use the arguments made in section.V to identify the geometry of the conformal manifolds, which is nothing but the coset space $G/H$ (see section.VIIB).
Since our primary goal is to find a general solution for the conformal manifolds in the parameter space, we use our understanding of the manifold geometry to develop a triad basis for the interaction matrix (section.VIIC). We then plug it back into the RGE in Eqn.27 to solve for the fixed points/manifolds. Once we solve for the conformal manifolds, we then study their stability properties (section.VIID). We then determine the phase boundary of the gapless surface(section.VIID). Finally, we show that we could study the infrared physics of the phase boundary via an effective Yukawa field theory wherein a single boson strongly interacts with $three$ flavors of surface fermions (VIIE).
VII.1 RG flow of the eigenvalues
Define $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ as the eigenvalues of the $\hat{V}$ matrix. Plugging the diagonal form of $\hat{V}$ into the RG equation in Eqn.27, we get the following decoupled equations for the eigenvalues,
@@ VII.1 RG flow of the eigenvalues
Table 3: Table showing the list of representative fixed points in terms of the eigenvalues of the real symmetric $\hat{V}$ matrix in the $\mathcal{N}=3$ TI surface (Eqn.50). Also shown is the shape of the manifold/coset space formed by each class of the fixed points in the 6-dimensional parameter space. (The fixed points with ${}^{\prime}*^{\prime}$ are representative points on the respective conformal manifolds)
@@
VII.2 Conformal manifolds as coset space $G/H$
Having determined the diagonal forms of the fixed points, we will now identify the manifold for each class of fixed points listed in Table 3. We repeat the procedure used for the $\mathcal{N}=2$ flavor surface. There, we identified the invariant subgroup $H$ of the fixed point matrix (see section.VIC). Then, the conformal manifold is the coset space $G/H$, where $G$ is the RGE’s symmetry. This method enables us to understand the geometry of all the conformal manifolds in the parameter space.

### v3

@@ V.4 Symmetries of the interaction matrix $V_{ij}$
V.5 Renormalization group study of the effective interacting model for a general $\mathcal{N}$: $2+\epsilon$ expansion
@@
So far, we have developed an effective interacting model for the TI surface. We are now ready to study the surface quantum criticality using the renormalization-group-equation (RGE) technique. In Eqn.LABEL:Lfrfmn, we have an interacting theory of $\mathcal{N}$ flavors of 2-component Dirac fermions in $(2+1)D$. Dimensional analysis suggests that the interaction term in Eqn.LABEL:Lfrfmn is marginal in $1+1$D dimensions. So in $(2+1)D$, the interaction term is irrelevant in the infrared (IR) limit if the interaction is weak. However, in the case of $\mathcal{N}=1$, it had been demonstrated that the interacting theory has an interacting UV fixed point [65, 66, 15, 67, 16, 17, 18]. Here, we study surface quantum criticality phenomena in a more general case with $\mathcal{N}>1$ surface Dirac cones and search for strong coupling conformal-field-theory fixed points that shall define possible universality classes of surface topological quantum criticality for a given $\mathcal{N}$.
@@
where $\epsilon=D-2$ and $D$ is the spacetime dimensions. Note that by increasing $l$, we are moving towards the IR limit of the theory. The numerical constant $c^{2}=\frac{\pi}{8}$. $V_{ij}$ is the renormalized dimensionless interaction matrix element. The derivation details of Eqn.27 are moved to the appendix.C for brevity.
@@
Note that the $SO(\mathcal{N})$ symmetry of the RG equation in Eqn.28 is $emergent$ and not fundamental. The theory does not have flavor rotation symmetry. The RG equation has this symmetry because we limited our calculations to one-loop contributions. A higher-loop contribution explicitly breaks the flavor rotation symmetry, as evident from the two- or three-loop diagrams shown in Fig.6 (b), (c), and (d).
Table 2: Table showing the list of fixed points, the shape of the manifold in the parameter space, and their stabilities for the $\mathcal{N}=2$ flavor surface. Here $\phi\in[0,2\pi]$ and $\epsilon=D-2$, where $D$ stands for spacetime dimensions.
@@
$\hat{V}_{c}=\frac{\epsilon}{2}\hat{I}+\frac{\epsilon}{2}\cos\phi\,\sigma_{z}+\frac{\epsilon}{2}\sin\phi\,\sigma_{x}$ One relevant, one irrelevant and one marginal operator. $\hat{I}$ ring
@@
VI Symmetries of the RGE and its dynamical solutions
Before we begin discussing the fixed points and their properties for different $\mathcal{N}$, here we present a fascinating observation from our study on the fixed points and the formation of conformal manifolds. Whether a set of fixed points forms a smooth manifold is conveniently and intuitively related to the dynamical symmetry of the RG equation and the invariant subgroup of the fixed point coupling matrix. In this section, we shall put forward the general argument. In the upcoming sections, where we solve for the fixed points of the RG Eqn.28 for $\mathcal{N}=2$ and $\mathcal{N}=3$ flavor surfaces, we shall show explicitly that this general argument is followed in all cases.
@@ VI Symmetries of the RGE and its dynamical solutions
If the dimension of the coset space is greater than zero, a smooth conformal manifold is formed in the parameter space. This symmetry argument proves to be very useful in understanding the shape and dimensions of the conformal manifolds if it exists. We shall make explicit use of this theory to solve the fixed points of the RG equation (Eqn.28) in the case of $\mathcal{N}=3$ TI surface, where the parameter space is six-dimensional and a direct brute-force calculation is challenging.
Interestingly, we shall find in section.IX that the symmetry of the dynamical equation is $not$ a necessary condition for the existence of conformal manifolds. When considering the 3-loop effects in the RG equation, we find that the $SO(\mathcal{N})$ symmetry is broken (see Fig.6). However, the conformal manifold persists, although in a distorted shape.
VII Conformal manifolds and phase boundary in the $\mathcal{N}=2$ flavor TI surface
@@
The main objective here is to study the isolated fixed points and conformal manifolds of the interacting theory in the three-dimensional parameter space ($D_{p}=3$). Once the fixed point manifolds are found, we intend to demonstrate how the conformal manifolds or isolated fixed points can be represented as a coset space $G/H$, as outlined in section.VI. Following this, we shall look for the phase boundary that separates the gapless surface state from the superconducting state. Finally, we show that the infrared physics at the phase boundary can be conveniently studied via an effective Yukawa field theory that involves a single complex boson strongly interacting with two flavors of gapless fermions.
The interaction matrix $\hat{V}$ defined in Eqn.28 is a $2\times 2$ symmetric matrix. Thus, the RG equation possesses emergent $SO(2)$ symmetry. Given that $\hat{V}$ is a real symmetric matrix, we could re-parametrize the $\hat{V}$-matrix in the following simple way:
@@
where $\sigma_{x}$ and $\sigma_{z}$ are Pauli matrices defined in the parameter space. This decomposition makes sense because a real symmetric matrix of size $2$ has $3$ independent parameters, which is the case here. Physically, $V_{x}$ signifies the scattering of Cooper pairs between the surface fermions of different flavors. $V_{z}$ represents the difference in the intra-flavor interactions of surface fermions. Plugging this back to the RG equation in Eqn.28, we get,
@@
Following the general discussions about the $SO(\mathcal{N})$ symmetry of Eq.28, 29, one can easily show that the RG flow lines exhibit an axial $SO(2)$ rotation symmetry. This is explicitly evident from the invariance of the RGE(given in Eqn.33) under $\hat{R}(\chi)$, $SO(2)$ rotations about the $V_{0}$ axis in the $V_{0}-V_{x}-V_{y}$ representation introduced in Eq.32,
@@
Another important consequence of the $SO(2)$ invariance is that the scaling dimensions of the operators are the same for all points on the ring manifold (see Appendix.D).
VII.1 Isolated fixed points Vs Conformal manifolds
@@
Figure 7: RG flows of the parameters $V_{0}$, $V_{z}$ and $V_{x}$ defined in Eqn.32. The red points denote the isolated fixed points, while the red-colored ring denotes the fixed point manifold. The blue line indicates the direction of an irrelevant operator corresponding to the $V_{x,c}=0,V_{z,c}=\frac{\epsilon}{2},V_{0,c}=\frac{\epsilon}{2}$ point on the ring manifold while the green line represents a relevant operator.
@@
$\hat{V}_{c}=\frac{\epsilon}{2}\hat{I}+\frac{\epsilon}{2}\cos\phi\,\sigma_{z}+\frac{\epsilon}{2}\sin\phi\,\sigma_{x},\,\phi\in[0,2\pi]$ (35)
@@
As discussed in section.III, a conformal manifold is associated with the presence of exactly marginal operators. We find one exact marginal operator at any point on the ring manifold, as expected since the manifold is one-dimensional. The marginal operator at a point identified by the metric $\phi$ on the manifold is given by,
$\mathcal{O}^{\phi}_{M}(x)$ $=$ $-\sin\phi\left(\Phi^{\dagger}_{1}\Phi_{1}(x)-\Phi^{\dagger}_{2}\Phi_{2}(x)\right)$ (37)
$+$ $\cos\phi\left(\Phi^{\dagger}_{1}\Phi_{2}(x)+\Phi^{\dagger}_{2}\Phi_{1}(x)\right)$
where $\Phi_{i}=\psi^{T}_{i}s_{y}\psi_{i}$ is the Cooper pair annihilation operator associated with the flavor ’$i$’($i=1,2$). Notice that the marginal operator is a function of the angle $\phi$, meaning that it is different at different locations on the manifold. It is directed tangential to the ring. Also the scaling dimension $\triangle_{\mathcal{O}_{M}}=D$, the spacetime dimensions.
We also find one relevant and one irrelevant operator on the manifold. Both the operators are functions of $\phi$. However, their scaling dimensions are independent of the location on the manifold, implying that the entire ring manifold belongs to a single universality class. See appendix.D for the exact form of the operators and their respective scaling dimensions.
The RG flow diagram is shown in fig.7. This ring manifold can be described as a conformal manifold of co-dimension $D_{co}=2$ in the three-dimensional parameter space.
VII.2 Conformal manifolds as a coset space $G/H$
Having found out all the fixed points from the RG equation, let us now see how the argument made in section VI on the connection between the dynamical symmetries and the geometry of the coset manifolds can be applied to each of the cases. We show here explicitly that the geometry of coset space for each fixed point matches what we found above using brute calculations.
@@ VII.3 Phase boundary
VII.4 Effective Yukawa field theory for the phase boundary
@@
Let us revisit the expression of the interaction matrix $\hat{V}_{c}$ at the conformal manifold once again. From the table.2, it has the form, $\hat{V}_{c}=\frac{\epsilon}{2}\hat{I}+\frac{\epsilon}{2}\cos\phi\,\sigma_{z}+\frac{\epsilon}{2}\sin\phi\,\sigma_{x}$. After simple rearrangements, $\hat{V}_{c}$ can be written as a tensor product of a vector in 2-dimensional space with itself. Let $\textbf{n}=\sqrt{\epsilon}\left(\cos\theta,\,\,\,\sin\theta\right)$ where $\theta=\phi/2$ be the unit vector. Then, $\hat{V}_{c}$ can be expressed in terms of n as,
@@
where $m$ is the boson mass and $\lambda$ is the four-boson interaction term. Here, we introduced a new term $g_{i}$ instead of $n_{i}$ used in Eqn.51 because after adding the boson dynamics into the theory, the effective Yukawa theory in Eqn.52 is quantitatively different from the one in Eqn.51. As mentioned at the beginning of this subsection, the boson mass term can be considered effectively as that relevant interaction that must be tuned to zero to study the quantum critical properties of the theory.
@@
VIII Conformal manifolds and phase boundary in the $\mathcal{N}=3$ flavor TI surface
@@
We constructed the effective 2D action governing the interaction in Eqns.22, LABEL:Lfrfmn. In section VD, we derived the RGE by studying the interaction up to the one-loop level. The interaction matrix $\hat{V}$ in Eqn.28 is now a $3\times 3$ symmetric matrix, and the interaction parameter space is six-dimensional ($D_{p}=6$). Therefore, a brute-force solution would be challenging. Here is where the emergent $SO(3)$ symmetry of the RGE comes to the rescue. It suggests that if they exist, the fixed point matrices $\hat{V}_{c}$ could be distinguished by their eigenvalues(or traces). Therefore, we shall first look at the RG flow of the eigenvalues of the $\hat{V}$ matrix in section.VIIIA.
Once we find the fixed points in terms of the eigenvalues, we use the arguments made in section.VI to identify the geometry of the conformal manifolds, which is nothing but the coset space $G/H$ (see section.VIIIB).
Since our primary goal is to find a general solution for the conformal manifolds in the parameter space, we use our understanding of the manifold geometry to develop a triad basis for the interaction matrix (section.VIIIC). We then plug it back into the RGE in Eqn.28 to solve for the fixed points/manifolds. Once we solve for the conformal manifolds, we then study their stability properties (section.VIIID). We then determine the phase boundary of the gapless surface(section.VIIID). Finally, we show that we could study the infrared physics of the phase boundary via an effective Yukawa field theory wherein a single boson strongly interacts with $three$ flavors of surface fermions (VIIIE).
VIII.1 RG flow of the eigenvalues
Define $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ as the eigenvalues of the $\hat{V}$ matrix. Plugging the diagonal form of $\hat{V}$ into the RG equation in Eqn.28, we get the following decoupled equations for the eigenvalues,
@@ VIII.1 RG flow of the eigenvalues
Table 3: Table showing the list of representative fixed points in terms of the eigenvalues of the real symmetric $\hat{V}$ matrix in the $\mathcal{N}=3$ TI surface (Eqn.53). Also shown is the shape of the manifold/coset space formed by each class of the fixed points in the 6-dimensional parameter space. (The fixed points with ${}^{\prime}*^{\prime}$ are representative points on the respective conformal manifolds). $\epsilon=D-2$, where $D$ stands for spacetime dimensions
@@
VIII.2 Conformal manifolds as coset space $G/H$
Having determined the diagonal forms of the fixed points, we will now identify the manifold for each class of fixed points listed in Table 3. We repeat the procedure used for the $\mathcal{N}=2$ flavor surface. There, we identified the invariant subgroup $H$ of the fixed point matrix (see section.VIIC). Then, the conformal manifold is the coset space $G/H$, where $G$ is the RGE’s symmetry. This method enables us to understand the geometry of all the conformal manifolds in the parameter space.